OpenStudy (anonymous):

medal will be given Explain how slope is used to determine if a triangle is a right triangle?

OpenStudy (anonymous):



OpenStudy (campbell_st):

well 2 lines are perpendicular if the product of there slopes is -1 perpendicular = right angle

OpenStudy (anonymous):

let m1 = the slope of AB. let m2 = slope of BC. the relation between two orthogonal lines is that m1 = -1/m2. if m1 = -1/m2, then it is a right triangle

OpenStudy (anonymous):

@urb0b thank you!